(b). A quality control analyst mixes 30.0 mL of 0.255 M BaCl, with excess AgNO, causing AgCl to precipitate and subsequently isolated 2.01 g of AgCl, Show a balanced equation for the reaction and calculate the percent yield of the reaction. (BaCl,: 208.23 g/mol. AgCt: 143.32 g/mol).
1
Expert's answer
2021-08-25T06:37:38-0400

AgNO 3 (aq) + BaCl 2 (aq) → 2 AgCl (s) + Ba(NO 3) 2 (aq)

Molar mass of BaCl2 = 208.23 g/mol

0.255× 208.23 = 53.09865

53.09865/30 = 1.77

= 1.77/2.01× 100

=88.06%